OpenStudy (anonymous):

how do you solve 3x^2+40=-x^2-56

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (anonymous):

3x^2+40= -x^2-56 3x^2 +x^2 = -56-40 4x^2 =-96 x^2 = -96/4 x^2 = -24 x=sq. root of -24 x =2sq.root of -6

OpenStudy (anonymous):

x= 2i sq.root of 6 where i = sq root of -1